OpenStudy (anonymous):

Suppose that the population of deer in a state is 1,500 and is growing 2% each year. Predict the population after 4 years.

OpenStudy (jackellyn):

1500(1 + 0.02)^4 = what?

OpenStudy (anonymous):

Use the formula: \[\large \color{green}{New \; Population = P(1 + \frac{R}{100})^t}\] Where P = Old Population, R = Rate of Growing and t = time...

OpenStudy (anonymous):

\[\huge \color{green}{\textbf{Welcome To Openstudy..}}\]

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Welcome dear..